OpenStudy (anonymous):

a body is thrown vertically up with a speed of 10 m/s from the top of a wall of height 3.45m the height from which another body should be dropped simultaneously such that two bodies reach the ground in same time (g=10m/s2) explanation??

OpenStudy (anonymous):

I don't understand why you are using 10m/s^2 for gravity. It should be 9.8

OpenStudy (anonymous):

IT IS THERE IN QUESTION (G=10M/S2

OpenStudy (anonymous):

that dosent really matter much romero infact it simplifies the question though i still havent figured it out :P

OpenStudy (anonymous):

Well it's not realistic. using 9.8 vs using 10 is a BIG difference that amounts to a great error.

OpenStudy (anonymous):

Ok so you have a wall that is 3.45 m right?

OpenStudy (anonymous):

Are both bodies released at the same time?

OpenStudy (anonymous):

well if both bodies start falling at the same time then all objects weather light or heavy reach the earth at same time as it dosent depend on mass

OpenStudy (anonymous):

it does nt matter about the mass

OpenStudy (anonymous):

options are 40m 24.5m 6.9m 20m and the answer is 24.5 but i dont know the steps to do it

OpenStudy (anonymous):

Oh easy just calculate the distance the first body travels vertically up before the force of gravity stops. AND the mass does matter because you need it to determine how much you need over time to make it stop.

OpenStudy (anonymous):

can u give the steps plzz

OpenStudy (anonymous):

I can't I don't know what the mass is.

OpenStudy (anonymous):

Actually you don't.

OpenStudy (anonymous):

Do you know the kinematic equations?

OpenStudy (anonymous):

no it dosent matter according s=vt or t=s/v the equation for acceleration due to gravity would be S=vf_vi+at where vf is the final speed vi is the initial speed

OpenStudy (anonymous):

I know I just want to know if the asker knew the what the kinematic equations were. - _ -

OpenStudy (anonymous):

yes i know the equations

OpenStudy (anonymous):

plz solve it for me.....

OpenStudy (anonymous):

shameer have uou written the complete numerical because it dosent say what should be calculared if you left something out please include

OpenStudy (anonymous):

*calculated

OpenStudy (anonymous):

the question is complete one i rechecked it...

OpenStudy (anonymous):

Well the kinematic equations involved 5 unknowns (actually 6 but we join to of them) t= time vi= initial velocity vf=final velocity a= acceleration d= distance traveled. How many of these do we know?

OpenStudy (anonymous):

acceleration and initial velocity are given

OpenStudy (anonymous):

how about vf=vi+gt

OpenStudy (anonymous):

we need to find the distance

OpenStudy (anonymous):

i am a little confused about the question when is the second body thrown downwards?

OpenStudy (anonymous):

two bodies are thrown and dropped at the same time

OpenStudy (anonymous):

@warz I understand the questions and the answer please let me finish and you can follow along if you want to understand it as well.

OpenStudy (anonymous):

vf2_vi2=2gS

OpenStudy (anonymous):

That would giv u the distance

OpenStudy (anonymous):

d= (vi + vf)/2 * t this also gives you the distrance.

OpenStudy (anonymous):

romero can u give the solved steps and answer plzzz

OpenStudy (anonymous):

Great so we know the initial velocity and the acceleration and we want the distance traveled. Like I said before the force of gravity will bring the body to stop at some point so actually know the final velocity.

OpenStudy (anonymous):

yes V=0

OpenStudy (anonymous):

with vi vf a and solving for d we use \[vf^2= vi^2 + 2*a*d\]

OpenStudy (anonymous):

Solve for d

OpenStudy (anonymous):

\[(vf^2 - vi^2)/2a= d\]

OpenStudy (anonymous):

plug in the given values vi=10m/s vf=0m/s a=10m/s^2

OpenStudy (anonymous):

then d= 5m

OpenStudy (anonymous):

yes note also that we have a negative vi but because gravity acts against the direct of the object acceleration will also be negative so that cancels the first negative and it makes it positive.

OpenStudy (anonymous):

ok then what to do???

OpenStudy (anonymous):

the final answer should be 24.5m

OpenStudy (anonymous):

Add it to the height of the cliff.

OpenStudy (anonymous):

so d=8.45 m

OpenStudy (anonymous):

next what to do??

OpenStudy (anonymous):

That should be your answer.

OpenStudy (anonymous):

but thats not the answer

OpenStudy (anonymous):

it is 24.5m but i dont know how to solve

OpenStudy (anonymous):

Then you wrote the question wrong. It's worded poorly and it might have confused me. I can see why I might have gotten the answer wrong. @warz also questioned if you wrote the question properly so you might want to go back and see if you wrote it right.

OpenStudy (anonymous):

a body is thrown vertically up with a speed of 10 m/s from the top of a wall of height 3.45m the height from which another body should be dropped simultaneously such that two bodies reach the ground in same time (g=10m/s2)

OpenStudy (anonymous):

i checked the question once more it does nt have any mistake...

OpenStudy (anonymous):

That "sentence" has a lot of grammar mistakes.

OpenStudy (anonymous):

it is the question from my book..

OpenStudy (aravindg):

any help needed?

OpenStudy (aravindg):

or am i late?

OpenStudy (anonymous):

he is right the question demands the explanation to how both bodies hit the ground at the same tim

OpenStudy (anonymous):

aravind help us out

OpenStudy (anonymous):

u r in at right time

OpenStudy (anonymous):

np you are not late you could help by explaining the question to us

OpenStudy (anonymous):

This can mean a lot of things. It can mean when the velocity of the body thrown up goes to zero and at the same time we have the second body dropped at the same time at the same location vf=0

OpenStudy (anonymous):

That's what I solved for.

OpenStudy (aravindg):

:) well basically the qn says that a 2 bodies 1-dropeed 2-thrown up they have same final velocities we need to find height from which second body was thrown

OpenStudy (aravindg):

lemme wrk out in a paper then i will giv the answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We know what it is :3.45m

OpenStudy (anonymous):

@AravindG can you explain me one thing about the question please is the second body thrown downwards at the same time the first is thrown upwards

OpenStudy (aravindg):

yes @warz

OpenStudy (aravindg):

ya i gt it!!

OpenStudy (anonymous):

The equation of kinematics for motion with constant acceleration in current case is \[h = h _{0}+v _{0}t-(g \times t^{2})/2\] [1] we put - sign before \[(g \times t^{2})/2\] because in our problem we have chosen upwards direction as positive, so acceleration due to gravity is negative because it Earth pulls down all objects on it. First body travels firstly upwards then downwards so in order to find entire time we must solve two equations: 1) v0=gt => t = v0/g or t1=10/10= 1s for upwards motion from first equation we find \[h = 2.2+10t-(10 \times 1^{2})/2=7.2 m\] it is the maximum height first object reaches 2) \[h=(g \times t ^{2}) /2\] => \[t=\sqrt{( 2 \times h)/g}=\sqrt{( 2 \times 7.2)/10}=1.2 s\] so \[t _{total}=1s+1.2s=2.2\] for first object. We apply the same procedure for second object and get 1) v0=gt => t = v0/g or t1=10/10= 1s for upwards motion from first equation we find \[h = 1.05+10t-(10 \times 1^{2})/2=6.05 m\] it is the maximum height second object reaches 2) \[h=(g \times t ^{2}) /2\] => \[t=\sqrt{( 2 \times h)/g}=\sqrt{( 2 \times 6.05)/10}=1.1 s\] so \[t _{total}=1s+1.1s=2.2\] for first object. Now the ratio of appropriate times is t1/t2=2.2/2.1=22/21 the answer is A.

OpenStudy (anonymous):

That's what I asked at first and you never said yes.

OpenStudy (aravindg):

A?

OpenStudy (anonymous):

Oh sorry for the second object \[t _{total}=1s+1.1s=2.1s\]

OpenStudy (aravindg):

@shameer1 are u satisfied with that answer ??else i will show my work

OpenStudy (anonymous):

@AravindG Let me get this straight. We know the first object is dropped at 3.45m . We want to know the height the second object is thrown up so that both objects hit the ground at the same time right?

OpenStudy (anonymous):

hahaha shameer u have got eve one scratching there heads i always used to keep a hand book for such buggers :D

OpenStudy (aravindg):

@Romero the second object is dropped down

OpenStudy (aravindg):

@shameer1 ,pls use use punctuations wherever necessary else the reader might get the qn wrong

OpenStudy (anonymous):

@AravindG We know the first object is thrown at 3.45m . We want to know the height the second object is dropped so that both objects hit the ground at the same time right? Better?

OpenStudy (anonymous):

hey aravind plz show ur work

OpenStudy (aravindg):

ya exactly

OpenStudy (aravindg):

@Romero u should include the word simultaneously

OpenStudy (anonymous):

the answer must be 24.5m

OpenStudy (aravindg):

k i will show my work

OpenStudy (anonymous):

the second object would be a chicken feather i guess thats why it floated cause of air friction :D

OpenStudy (anonymous):

@Shameer have you changed the text of question?

OpenStudy (anonymous):

see no one understood what the question asked for. lol worded poorly.

OpenStudy (anonymous):

hey that was just such a good explanation that a body thrown fa greater hight would have greater speed

OpenStudy (anonymous):

vf2_vi2=2gh greater the value of h greater the speed bingo

OpenStudy (aravindg):

well first we calculate time taken by the first body , consider the upward motion of it well the time taken can be found out using \[\large v=u+at\] \[\large 0=10-10t\] which gives t=1s the distance travelled upward will be \[H=10t-\frac{1}{2}\times5\times1^2\] \[=5m\] now so at topmost point height of first body will be \[\large 3.45+5=8.45\] \[\large8.45=\frac{1}{2}\times10\times t^2\] solving which we get t=1.3s

OpenStudy (aravindg):

now for both bodies the time is same 1.3s

OpenStudy (aravindg):

wait a sec the height of second body i am getting is 8.45m

OpenStudy (anonymous):

both bodies actually travel 2.3 s you forgot about the 1s

OpenStudy (aravindg):

oops ths romero

OpenStudy (aravindg):

thx

OpenStudy (anonymous):

@Romero exactly

OpenStudy (anonymous):

we need to find distance

OpenStudy (aravindg):

ya i gt 26.45